# Permutation-Invariant Spectral Learning via Dyson Diffusion

### v1

@@ Appendix H Inference
6:  $\delta t\leftarrow\text{BackwardStepsizeController}(\lambda(t),u)$ $\triangleright$ Step size as described F.2.
@@
8:  $\rm BackwardsDrift$$\leftarrow$ drift of eq. 5 at time $t$ and point $\lambda(t)$ with score $s_{\theta}(\lambda,t_{j+1})$
@@
Figure 6: Sampling from the Dyson Diffusion Model. The eq. 5 is evolved backward in time with the shooting mechanism and an adaptive step size, ensuring that the paths remain in the Weyl Chamber.
We describe in Algorithm 3 the sampling procedure with the shooting mechanism (see Appendix G for details) incorporated, ensuring that we remain in the Weyl Chamber.
Appendix I Engineering
We follow Karras et al. (2024) by using SiLU activations. We further use EMA to average over multiple runs (Song & Ermon, 2020).
A key strength of DyDM is that the score $s_{\theta}(\lambda,t)$can be parameterized with any learning architecture, without being constrained to GNNs or graph transformers. We demonstrate this by parameterizing the score with a simple MLP. The size of layers varies by application, and we document it for each dataset in Github . For instance, for the large dataset of $15^{\prime}000$ brain ego graphs, our model consists of a hidden MLP of depth $4$, where the input and output layer have width $64$ and the hidden layers have width $256$. The space + time data is first scaled up with a linear layer from size $n+1=11$ followed by a batch norm to feed into the hidden MLP, and upon processing through the hidden MLP, it gets scaled down through a simple linear layer to size $n=10$. In the hidden MLP, we employ as nonlinearities scaled SiLU functions, as argued by Karras et al. (2024).
@@ Appendix I Engineering
We implement the model in Jax and Equinox (Kidger & Garcia, 2021).
I.1 Choice of Time Grid
As outlined in Section 3.2, we choose an exponential time grid on which the objective is learned. This is due to the mixing behavior of Dyson’s Brownian Motion. For instance, for the Brain dataset, we use exponential time grid detailed in Table 2.
Table 2: Example of exponential time grid, here for the brain data set which contains in total $15^{\prime}000$ graphs. $\mathrm{d}t$ is $0.05$, and the final time is $T=12.0$.
@@ 77 T $2\cdot\mathrm{d}t$
I.2 Preprocessing and dealing with different dimensions
As is the case for an Ornstein-Uhlenbeck diffusion, the speed of convergence depends for Dyson-BM on $(1)$ the coefficients and $(2)$ the initial condition. We can thus choose to significantly vary the final time $T$ or rescale the initial condition. We choose the latter (although both options are feasible). To that end, we rescale the spectra with an affine transformation in a preprocessing step. In more detail: For a given set of graphs, we perform an eigendecomposition, and rescale so that among the the entire dataset the largest eigenvalue is at most $\lambda_{\rm max}$ and at least $\lambda_{\rm min}$. For instance, on the benchmark models we chose $\lambda_{\rm max}=5$, $\lambda_{\rm min}=-5$. If a spectrum has eigenvalues of multiplicities greater than $1$, we perform an $\epsilon$-perturbation, where the $\epsilon$ depends on the distance of the closest eigenvalues in the dataset. In postprocessing, the spectra are scaled back and the perturbation is undone for eigenvalues that are at most $\epsilon$ apart after generation. Note that with this preprocessing, only one eigendecomposition per training graph is necessary.
Appendix J Challenges of Dyson’s Brownian Motion
@@ I.2 Preprocessing and dealing with different dimensions
We overcome the obstacles mentioned above and provide a diffusion model for the spectra of graphs based on the Dyson SDE (Fig. 1). The model is not only efficient but is also able to distinguish between spectra of graphs that GNNs are blind to (Fig. 2). In addition, with DyDM, no ad hoc data augmentation is necessary. Further, through Eigenvector-SDE, we give the dynamics of the remaining information in form of conditioned eigenvector dynamics, hence making them accessible for future work devoted to eigenvector diffusion.
J.1 Why not log-transform the SDE?
One idea could be to transform the spectral SDE into terms $\lambda_{1},\lambda_{2}-\lambda_{3},\ldots,\lambda_{n-1}-\lambda_{n}$ and take logarithms, to avoid singularities. However, this is no desirable for multiple reasons: First, upon applying Ito to this SDE, we (i) lose the $\log$ and obtain singularities again and (ii) get higher order singularities $\mathrm{d}\log(x_{t})=\frac{1}{x_{t}}\mathrm{d}x_{t}-\frac{1}{2x_{t}^{2}}(\mathrm{d}x_{t})^{2}$. Second, the space that would need to be sampled would certainly not decrease, since now the transformed domain reaches from -$\infty$ to $\infty$. Hence, we choose the method described in the main part.
Appendix K Why not learn on all $n!$ many graph representations?
In short, learning $n!$ more data is much harder. This point has been mentioned by previous literature, e.g. Niu et al. (2020). However, if we go deeper, the interested reader might wonder why exactly.
K.1 Rigorous argument
@@
Corollary K.1 (Mean square error).
@@
To prove Corollary K.1, we derive the mean squared error for both estimators.
Lemma K.2 (MSE of $\rm est_{A}$}).
@@
By considering the estimation problem as a problem of parameters $m$, NN, Corollary K.1 follows directly from Lemma K.2, Lemma K.3.
@@
Appendix L WL-equivalence of regular graphs
We now prove Lemma 2.1.
Lemma 2.1 (restated).
@@
For a $k$-regular graph, since every vertex has exactly $k$-neighbors, in each iteration of the WL-algorithm all vertices have the same color. Thus in particular the histograms are always the same and therefore any two $k$-regular graphs are WL-equivalent.
A $k$-regular graph is not WL-equivalent to a $\ell$-regular graph for $k\neq\ell$ since the colors after the first iteration are distinct as the number of neighbors is different. Moreover, if a graph is not regular, then the first iteration must assign a different new color to at least two vertices in the first iteration. Therefore, the color histogram is not the same as the color histogram of a regular graph and thus they are not WL-equivalent. This shows that the WL-equivalence class of a $k$-regular graph is exactly the set $k$-regular graphs. ∎
In particular, it follows from Morris (Morris et al., 2019) that GNNs cannot distinguish $k$-regular graphs.
Appendix M Datasets
@@
The WL-Bimodal graph consists of $80\%$ graph $A$ and $20\%$ graph B (see Fig. 2) adjacency matrices. We drew among all permutations $5^{\prime}000$ permutations uniformly at random and shuffled the graphs. The first $80\%$ of this dataset are used for training, the remaining $20\%$ are used for testing.
@@
This standard benchmark dataset (Niu et al., 2020; Jo et al., 2022; You et al., 2018) consists of $100$ graphs of size up to $20$ vertices. We comment in Section N.1 on the small dataset set ($100$ graphs) compared to the big dimension (up to 20 vertices) and the thereby induced effect of undersampling.
@@
We report this dataset in our repository. In detail, we construct from the brain graph Amunts et al. (2013); Rossi & Ahmed (2015) so-called ego-graphs. That is, we take the (distance 1) neighborhoods of vertices, and consider the induced subgraph. From those, we generate $15^{\prime}000$ graphs of size $n=5$ to $n=10$ vertices with eigenvector multiplicity up to $3$, with the closest eigenvalues – which are not multiplicities – having distance $0.036$. We take $70\%$ as train graphs, $15\%$ as validation, and the remaining $15\%$ as test graphs.
Appendix N Comparing to Benchmark Models
N.1 On undersampling
@@
For the “bimodal” case, we have sufficient statistics for the $10$-dimensional space $C_{10}$ ($N=5^{\prime}000$ graph samples, each isomorphic to one of two graphs) and know in addition the underlying distribution; hence, an extensive interpretation of this result is appropriate. For a fair comparison, we thus follow the standard test/train split procedure as reported in (Jo et al., 2022; You et al., 2018; Niu et al., 2020) using $80\%$ of the data as train data and the remaining $20\%$ as test data.
@@
Conversely, the standard benchmark set “community small” (Niu et al., 2020; Jo et al., 2022; You et al., 2018) contains only 100 graphs, and each has a size of up to $n=20$ vertices. Thus, a comparison from the learned distribution based on (few) training samples to (very few) test samples suffers from undersampling. This becomes very stark if one considers the following issue: If one would extract $80\%$ of the dataset for training, where they are taken from already matters a lot: Whether they are taken from the front or back changes the maximal graph size in the training set. Depending on whether the training data is taken from the front or back of the dataset, a model trained on the training set might thus have no possibility to learn the correct maximal graph size. More generally, poorness of benchmarks in graph generative learning has been recently addressed by Bechler-Speicher et al. (2025). To offer some consistent comparison, we do include the standard benchmark “community small”, but focus on memorization rather than the (on those benchmarks untestable) generalization.
@@
To overcome the issue of undersampling, we construct a set of $15^{\prime}000$ ego-graphs from the brain dataset Amunts et al. (2013) as described in Appendix M, which is sufficiently large to not suffer from undersampling. We train both our model and DiGress on $70\%$ (= $10^{\prime}500$ graphs), perform hyperparameter tuning (see below for details on the DiGress hyperparameter tuning) on a validation set of $15\%$, and test on the remaining $15\%$.
N.2 GDSS
@@
To compare to the GDSS model Jo et al. (2022), we take the following approach to ensure a fair comparison: The GDSS model has been trained and optimised on the ego small and community small dataset, so that we take for these datasets the snapshots and hyperparameters given by the original paper Jo et al. (2022). For the remaining datasets, we start with the settings from the community-small dataset since that has similar size, adapt the maximal number of vertices to the dataset, and then perform hyperparameter tuning as described in Appendix C of the GDSS paper: We form a grid search on the model’s following parameters: The scale coefficient in $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$, the signal-to-noise ratio in $\{0.05,0.1,0.15,0.2\}$, and in addition to the GDSS paper, we also try different $\beta_{max}$ in $\{1,10,20\}$, and try different batch sizes $\{128,4096\}$. We test with and without EMA. The motivation for the additional parameters we hyper tune is that we observed that they further improve the GDSS model. This gives full fairness to the model. The training time on one H100 GPU per job was about 24 hours for each of the configurations with batch size $128$ and approximately 1 hour and 45 minutes for each of the configurations with batch size $4096$.
@@
On the Two-WL graph case, our hyperparameter tuning resulted in $280$ models that we trained and sampled from. From each of those 280 models, we generated $1^{\prime}000$ graphs. Most of those models worked fine, that is 269 models did not contain NaNs in their output. From those, we select the best model based on the following relative error: From the generated samples, we calculate the share of spectra $\epsilon$-close to the spectrum of graph $A$, say $\hat{p}_{A}$ and the share of spectra $\epsilon$-close to graph $\hat{p}_{B}$ (we choose the $l_{2}$ distance with $\epsilon=0.2$). Recall that in the ground truth, we have $p_{A}=0.8$, $p_{B}=0.2$. We then selected the best model based on the relative error
@@
N.3 EDP-GNN
We proceeded analogously to GDSS: We used the given configurations for community small. We observe that the model already performs hyperparameter tuning of the noise scales during sampling. For the other data sets, we have used the given configurations for community small and in addition tried the learning rates $\{0.001,0.0002\}$, number of diffusion steps $\{1^{\prime}000,2^{\prime}000\}$ and number of layers $\{4,6\}$. In the two graph case, for example, the optimal configuration was with learning rate $0.0002$, number of diffusion steps being equal to $2^{\prime}000$ and $6$ layers. The training time for $5^{\prime}000$ epochs on a H100 GPU was approximately 10 hours.
N.4 DiGress
As with the previous models, we used the given configurations for community small. For the other data sets, we have used the given configurations for community small and in addition tried the learning rates $\{0.001,0.0002\}$, weight decay parameters $\{10^{-2},10^{-12}\}$, number of diffusion steps $\{500,1^{\prime}000\}$ and number of layers $\{5,8\}$. The model was quite robust to hyperparameter tuning and in the two graph example, for all parameters the model sampled the $80\%$-graph with likelihood between $75\%$ and $82\%$. The training time for $1^{\prime}000$ epochs on a H100 GPU was approximately 3 hours.
N.5 Congress
Since Congress and DiGress come from the same paper, we have proceeded almost exactly as in Digress. We tried the learning rates $\{0.001,0.0002\}$, weight decay parameters $\{10^{-2},10^{-12}\}$, number of diffusion steps $\{500,1^{\prime}000\}$ and number of layers $\{6,8\}$. The model was again rather robust to hyperparameter tuning, yet not as much as Digress. In the two graph example, for all parameters the model sampled the $80\%$-graph with likelihood between $10\%$ and $25\%$. The default learning rate from the community-small configuration was $0.0002$, yet we have observed that the results were significantly better with learning rate $0.001$. The training time for $1^{\prime}000$ epochs on a H100 GPU was approximately 3 hours.
N.6 Comparison table with more digits
We provide here the table shown in the main part Table 1 but with more digits (not implying that all are statistically significant): This rationalizes which entries in Table 1 are dark green and which ones are light green. Note that this result is included only for transparency, since the reported digits here go beyond the significant digits.
Table 3: Statistical distances of DyDM compared to standard models, as in Table 1 but with more digits.
Dataset
WL-Bimodal
Community Small Brain
@@
EDP-GNN 0.1342 0.0750 0.4164 0.1356 - -
GDSS 0.2289 0.1252 0.4180 0.1444 - -
ConGress 0.3802 0.1590 0.2741 0.1138 - -
DiGress (no trick) 1.0568 0.2852 2.5088 0.4481 - -
@@
Appendix O Learning dynamics of EDP-GNN
We report in Figure 7 the learning progress of EDP-GNN on the WL-bimodal dataset. We average over $4$ different training and sampling runs of EDP-GNN. After $5^{\prime}000$ epochs, we observe the result of EDP-GNN reported in Figure 2.
We observe that the model quickly learns the WL equivalence class (the pink line is from epoch $500$ onward close to $1$). The share of graph $A$ and graph $B$ samples initially increase until epoch $1^{\prime}500$, but then remain low and significantly different from the ground truth (blue and green dashed lines). Importantly, a significant share of the samples are $WL$-equivalent but neither isomorphic to graph $A$ nor to graph $B$.

### v2

@@ Appendix I Empirical Considerations
6:  $\delta t\leftarrow\text{BackwardStepsizeController}(\lambda(t),u)$ $\triangleright$ Step size as described in  F.2.
@@
8:  $\rm BackwardsDrift$$\leftarrow$ drift of Equation˜4 at time $t$ and point $\lambda(t)$ with score $\hat{s}_{\theta}(\lambda,t_{j+1})$
@@
Figure 7: Sampling from the Dyson Diffusion Model. The Equation˜4 is evolved backward in time with the shooting mechanism and an adaptive step size, ensuring that the paths remain in the Weyl Chamber.
We describe in Algorithm˜3 the sampling procedure with the shooting mechanism (see Appendix˜G for details) incorporated, ensuring that we remain in the Weyl Chamber.
Appendix K Engineering
We follow [50] by using SiLU activations. We further use EMA to average over multiple runs [51].
A key strength of DyDM is that the score $s_{\theta}(\lambda,t)$ can be parameterized with any learning architecture, without being constrained to GNNs or graph transformers. We demonstrate this by parameterizing the score with a simple MLP. The size of layers varies by application, and we document it for each dataset in the uploaded Supplementary Material. For instance, for the large dataset of $15,000$ brain ego graphs, our model consists of a hidden MLP of depth $4$, where the input and output layer have width $64$ and the hidden layers have width $256$. The space + time data is first scaled up with a linear layer from size $n+1=11$ followed by a batch norm to feed into the hidden MLP, and upon processing through the hidden MLP, it gets scaled down through a simple linear layer to size $n=10$. In the hidden MLP, we employ as nonlinearities scaled SiLU functions, as argued by [50]. We use AdamW for optimization [52].
@@
Our runtimes are as follows: For WL-Bimodal, sampling $1000$ paths takes $16$ seconds and training takes $19$ hours. For Community Small, sampling $1000$ paths takes $60$ seconds and training takes $10$ hours. For Brain, sampling of $1000$ paths is very fast at $11$ seconds. Training takes $16$ hours. Runtime-wise, our strength therefore lies in fast generation.
We sample $1000$ paths to make the reported distances statistically robust, and report in addition to the mean also the Wasserstein of the marginals of these $1000$ samples to the ground truth, accounting both for high-dimensional and marginal effects (see Section˜4.2).
We implement the model in JAX [53] (version 0.9.0, licensed under Apache license, version 2.0) and Equinox https://github.com/patrick-kidger/equinox [54] (version 0.13.6, licensed under Apache license, version 2.0).
K.1 Choice of Time Grid
As outlined in Section˜3.2, we choose an exponential time grid on which the objective is learned. This is due to the mixing behavior of Dyson’s Brownian motion. For instance, for the Brain dataset, we use an exponential time grid as detailed in Table˜2.
Table 2: Example of exponential time grid, here for the brain dataset which contains in total $15,000$ graphs. $\mathrm{d}t$ is $0.05$, and the final time is $T=12.0$.
@@ 77 T $2\cdot\mathrm{d}t$
K.2 Preprocessing and dealing with different dimensions
As is the case for an Ornstein-Uhlenbeck diffusion, the speed of convergence depends for (Dyson-BM) on $(1)$ the coefficients and $(2)$ the initial condition. We can thus choose to significantly vary the final time $T$ or rescale the initial condition. We choose the latter (although both options are feasible). To that end, we rescale the spectra with an affine transformation in a preprocessing step. In more detail: For a given set of graphs, we perform an eigendecomposition, and rescale so that among the entire dataset the largest eigenvalue is at most $\lambda_{\rm max}$ and at least $\lambda_{\rm min}$. For instance, on the benchmark models we chose $\lambda_{\rm max}=5$, $\lambda_{\rm min}=-5$. If a spectrum has eigenvalues of multiplicities greater than $1$, we perform an $\epsilon$-perturbation, where the $\epsilon$ depends on the distance of the closest eigenvalues in the dataset. In postprocessing, the spectra are scaled back and the perturbation is undone for eigenvalues that are at most $\epsilon$ apart after generation. Note that with this preprocessing, only one eigendecomposition per training graph is necessary.
Appendix L Challenges of Dyson’s Brownian Motion
@@
We overcome the obstacles mentioned above and provide a diffusion model for the spectra of graphs based on the Dyson SDE (Fig. 1). The model is not only efficient but is also able to distinguish between spectra of graphs that GNNs are blind to (Fig. 2). In addition, with DyDM, no ad hoc data augmentation is necessary. Further, through (Eigenvector-SDE), we give the dynamics of the remaining information in form of conditioned eigenvector dynamics, hence making them accessible for future work devoted to eigenvector diffusion.
L.1 Why not log-transform the SDE?
One idea could be to transform the spectral SDE into terms $\lambda_{1},\lambda_{2}-\lambda_{3},\ldots,\lambda_{n-1}-\lambda_{n}$ and take logarithms, to avoid singularities. However, this is not desirable for multiple reasons: First, upon applying Ito to this SDE, we (i) lose the $\ln$ and obtain singularities again and (ii) get higher order singularities $\mathrm{d}\ln(x_{t})=\frac{1}{x_{t}}\mathrm{d}x_{t}-\frac{1}{2x_{t}^{2}}(\mathrm{d}x_{t})^{2}$. Second, the space that would need to be sampled would certainly not decrease, since now the transformed domain reaches from -$\infty$ to $\infty$. Hence, we choose the method described in the main part.
Appendix M Why not learn on all $n!$ many graph representations?
In short, learning $n!$ more data is much harder. This point has been mentioned by previous literature, e.g. [8]. However, if we go deeper, the interested reader might wonder why exactly.
M.1 Rigorous argument
@@
Theorem 6 (Mean square error).
@@
To prove Theorem˜6, we derive the mean squared error for both estimators.
Lemma 7 (MSE of $\rm est_{A}$}).
@@
By considering the estimation problem as a problem of parameters $m$, NN, Theorem˜6 follows directly from Lemma˜7, Lemma˜8.
@@
Appendix N WL-equivalence of regular graphs
We now prove Lemma˜1.
Lemma 1 (restated).
@@
For a $k$-regular graph, since every vertex has exactly $k$ neighbors, in each iteration of the WL-algorithm all vertices have the same color. Thus in particular the histograms are always the same and therefore any two $k$-regular graphs are WL-equivalent.
A $k$-regular graph is not WL-equivalent to a $\ell$-regular graph for $k\neq\ell$ since the colors after the first iteration are distinct as the number of neighbors is different. Moreover, if a graph is not regular, then the first iteration must assign a different new color to at least two vertices in the first iteration. Therefore, the color histogram is not the same as the color histogram of a regular graph and thus they are not WL-equivalent. This shows that the WL-equivalence class of a $k$-regular graph is exactly the set of $k$-regular graphs. ∎
In particular, it follows from Morris [14] that GNNs cannot distinguish $k$-regular graphs.
Appendix O Laplacian Spectrum
The Laplacian spectrum captures key properties of a graph, such as the Fiedler Value encoding the algebraic connectivity of the graph. Our model works perfectly well on the Laplacian spectrum, for which we plot the marginal densities on the datasets “Brain” and “Community Small”. Note that the model perfectly learns key properties such as the fact that the smallest eigenvalue is 0 and learns the distributions well.
For Brain (Figure˜8), we plot for each eigenvalue $\lambda_{i}$ a histogram of the DyDM-generated marginal (blue) compared to a histogram of the Ground Truth (note that the Brain dataset test for generalization). For the commonly used “Community Small” dataset, we show in Figure˜9 the first $12$ eigenvalues, since only $73$ graphs have more than $12$ nodes.
Note that in this Section, we use the convention $\lambda_{1}\leq\cdots\leq\lambda_{n}$ to stay consistent with Graph Laplacian literature.
Figure 8: Performance of DyDM on the Laplacian of the Brain dataset compared to the underlying distribution: DyDM perfectly learns the (rough) distribution and key properties such as the smallest eigenvalue always being 0 (faint line, since it is very concentrated) or the distribution of the Fiedler Value ($\lambda_{2}$).
Figure 9: Performance of DyDM on the Laplacian of the Community Small dataset: DyDM learns the distribution and key properties very well.
Appendix P Datasets
@@
The WL-Bimodal dataset consists of $80\%$ graph $A$ and $20\%$ graph B (see Fig. 2) adjacency matrices. We drew among all permutations $5,000$ permutations uniformly at random and shuffled the graphs. Hence, we have $5,000$ matrices, of which $4,000$ represent graph $A$ and $1,000$ represent graph $B$. Adjacency matrices representing the same graph are not identical matrices, but rather permutations of each other, making this dataset relevant to graph generative models. The first $80\%$ of this dataset are used for training, the remaining $20\%$ are used for testing.
@@
This standard benchmark dataset [8, 9, 31] consists of $100$ graphs of size up to $20$ vertices. We comment in Section˜Q.1 on the small dataset ($100$ graphs) compared to the big dimension (up to 20 vertices) and the thereby induced effect of undersampling.
@@
We report this dataset in our repository. In detail, we construct from the brain graph [32, 33] so-called ego-graphs. That is, we take the (distance 1) neighborhoods of vertices, and consider the induced subgraph. From those, we generate $15,000$ graphs of size $n=5$ to $n=10$ vertices with eigenvalue multiplicity up to $3$, with the closest eigenvalues – which are not multiplicities – having distance $0.036$. We take $70\%$ as train graphs, $15\%$ as validation, and the remaining $15\%$ as test graphs.
Appendix Q Comparing to Benchmark Models
Q.1 On undersampling
@@
For the “bimodal” case, we have sufficient statistics for the $10$-dimensional space $C_{10}$ ($N=5,000$ graph samples, each isomorphic to one of two graphs) and know in addition the underlying distribution; hence, an extensive interpretation of this result is appropriate. For a fair comparison, we thus follow the standard test/train split procedure as reported in [9, 31, 8] using $80\%$ of the data as train data and the remaining $20\%$ as test data.
@@
Conversely, the standard benchmark set “community small” [8, 9, 31] contains only 100 graphs, and each has a size of up to $n=20$ vertices. Thus, a comparison from the learned distribution based on (few) training samples to (very few) test samples suffers from undersampling. This becomes very stark if one considers the following issue: If one would extract $80\%$ of the dataset for training, where they are taken from already matters a lot: Whether they are taken from the front or back changes the maximal graph size in the training set. Depending on whether the training data is taken from the front or back of the dataset, a model trained on the training set might thus have no possibility to learn the correct maximal graph size. More generally, poorness of benchmarks in graph generative learning has been recently addressed by [30]. To offer some consistent comparison, we do include the standard benchmark “community small”, but focus on memorization rather than the (on those benchmarks untestable) generalization.
@@
To overcome the issue of undersampling, we construct a set of $15,000$ ego-graphs from the brain dataset [32] as described in Appendix˜P, which is sufficiently large to not suffer from undersampling. We train both our model and DiGress on $70\%$ (= $10,500$ graphs), perform hyperparameter tuning (see below for details on the DiGress hyperparameter tuning) on a validation set of $15\%$, and test on the remaining $15\%$.
Q.2 GDSS
@@
To compare to the GDSS model [9], we take the following approach to ensure a fair comparison: The GDSS model has been trained and optimised on the ego small and community small dataset, so that we take for these datasets the snapshots and hyperparameters given by the original paper [9]. For the remaining datasets, we start with the settings from the community-small dataset since that has similar size, adapt the maximal number of vertices to the dataset, and then perform hyperparameter tuning as described in Appendix C of the GDSS paper: We form a grid search on the model’s following parameters: The scale coefficient in $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$, the signal-to-noise ratio in $\{0.05,0.1,0.15,0.2\}$, and in addition to the GDSS paper, we also try different $\beta_{max}$ in $\{1,10,20\}$, and try different batch sizes $\{128,4096\}$. We test with and without EMA. The motivation for the additional parameters we hyper tune is that we observed that they further improve the GDSS model. This gives full fairness to the model. The training time on one A100 GPU per job was for the brain dataset about 24 hours for each of the configurations with batch size $128$ and approximately 1 hour and 45 minutes for each of the configurations with batch size $4096$.
For the community small dataset, training took $15$ minutes, sampling $450$ seconds.
For the WL-Bimodal case, again using an $A100$ GPU, one training run of GDSS took $5$ hours, and one sampling run of 1,0001{,}000 graphs took $450$ seconds.
@@
On the Two-WL graph case, our hyperparameter tuning resulted in $280$ combinations that we trained and sampled from. From each of those 280 models, we generated $1,000$ graphs. Most of those models worked fine, that is, $269$ models did not contain NaNs in their output. From those, we select the best model based on the following relative error: From the generated samples, we calculate the share of spectra $\epsilon$-close to the spectrum of graph $A$, say $\hat{p}_{A}$ and the share of spectra $\epsilon$-close to graph $\hat{p}_{B}$ (we choose the $l_{2}$ distance with $\epsilon=0.2$). Recall that in the ground truth, we have $p_{A}=0.8$, $p_{B}=0.2$. We then selected the best model based on the relative error
@@
We then trained the best hyperparameter choice with $5$ different seed values and report their mean and standard deviation.
@@
Q.3 EDP-GNN
We proceeded analogously to GDSS: We used the given configurations for community small. We observe that the model already performs hyperparameter tuning of the noise scales during sampling. For the other datasets, we have used the given configurations for community small and in addition tried the learning rates $\{0.001,0.0002\}$, number of diffusion steps $\{1,000,2,000\}$ and number of layers $\{4,6\}$. In the two graph case, for example, the optimal configuration was with learning rate $0.0002$, number of diffusion steps being equal to $2,000$ and $6$ layers. The training time for $5,000$ epochs on a H100 GPU was approximately 10 hours.
Q.4 DiGress
As with the previous models, we used the given configurations for community small. For the other datasets, we have used the given configurations for community small and in addition tried the learning rates $\{0.001,0.0002\}$, weight decay parameters $\{10^{-2},10^{-12}\}$, number of diffusion steps $\{500,1,000\}$ and number of layers $\{5,8\}$. The model was quite robust to hyperparameter tuning and in the two graph example, for all parameters the model sampled the $80\%$-graph with likelihood between $75\%$ and $82\%$. The training time for $1,000$ epochs on a H100 GPU was approximately 3 hours.
Q.5 ConGress
Since ConGress and DiGress come from the same paper, we have proceeded almost exactly as in DiGress. We tried the learning rates $\{0.001,0.0002\}$, weight decay parameters $\{10^{-2},10^{-12}\}$, number of diffusion steps $\{500,1,000\}$ and number of layers $\{6,8\}$. The model was again rather robust to hyperparameter tuning, yet not as much as DiGress. In the two graph example, for all parameters the model sampled the $80\%$-graph with likelihood between $10\%$ and $25\%$. The default learning rate from the community-small configuration was $0.0002$, yet we have observed that the results were significantly better with learning rate $0.001$. The training time for $1,000$ epochs on a H100 GPU was approximately 3 hours.
Q.6 Comparison table with more digits
We provide here the table shown in the main part Table˜1 but with more digits (not implying that all are statistically significant): This rationalizes which entries in Table˜1 are dark green and which ones are light green. Note that this result is included only for transparency, since the reported digits here go beyond the significant digits.
Table 3: Statistical distances of DyDM compared to standard models, as in Table˜1 but with more digits.
Dataset WL-Bimodal Community Small Brain
@@
EDP-GNN 0.1342 0.0750 0.4164 0.1356 0.0723 0.0321
GDSS 0.2416 0.1339 0.3812 0.1444 0.3276 0.1191
ConGress 0.3802 0.1590 0.2741 0.1138 0.1315 0.0296
DiGress (no trick) 1.0568 0.2852 2.5088 0.4481 0.5686 0.1749
@@
Table 4: Empirical standard deviation of runs.
Dataset WL-Bimodal Community Small Brain
DyDM (ours) 0.0057 0.0011 0.0341 0.0044 - -
EDP-GNN 0.0251 0.0226 - - 0.0838 0.0265
GDSS 0.0125 0.0067 0.0617 0.0055 0.0951 0.0099
ConGress 0.3143 0.0736 0.0171 0.0033 0.0943 0.0131
DiGress (no trick) 0.0349 0.0060 0.0169 0.0057 0.4252 0.1293
DiGress (trick) 0.0150 0.0033 0.0129 0.0025 0.0458 0.0065
Appendix R Learning dynamics of EDP-GNN
We report in Figure˜10 the learning progress of EDP-GNN on the WL-Bimodal dataset. We average over $4$ different training and sampling runs of EDP-GNN. After $5,000$ epochs, we observe the result of EDP-GNN reported in Figure˜2.
We observe that the model quickly learns the WL equivalence class (the pink line is from epoch $500$ onward close to $1$). The share of graph $A$ and graph $B$ samples initially increase until epoch $1,500$, but then remain low and significantly different from the ground truth (blue and green dashed lines). Importantly, a significant share of the samples are WL-equivalent but neither isomorphic to graph $A$ nor to graph $B$.